OpenStudy (anonymous):

what is the slope of r=squareroot of cos^2 theta at theta= pi/4

OpenStudy (amistre64):

r = sqrt(cos^2(pi/4)) cos(45) = sqrt(2)/2 square that and we get 2/4 r = sqrt(2/4) r = sqrt(2)/2

OpenStudy (amistre64):

the slope? it is the slope of 45 degrees if I recall correctly. slope = 1